Speedys's Pizza produces fresh pizzas for area grocery stores that sell them for $10 each. Speedy's production cost have been estimated as
TC = 1,000 - 50Q + 5Q^2
(a) What is the firm's marginal cost function ?
(b) What output rate will maximize speedy's profits ?
1
Expert's answer
2018-03-12T06:41:07-0400
Answer:
a) MC=-50+10*Q
b) -50+10*Q=10
10*Q=60
Q=6